What mass of silver chromate is produced from 1.23 mol of silver nitrate
1
Expert's answer
2020-03-30T10:52:07-0400


What mass of silver chromate is produced from 1.23 mol of silver nitrate.


Solution:

The chemical reaction of silver chromate formation is following:

2AgNO3 + K2CrO4 = Ag2CrO4 (s) + 2KNO3

We can calculate the mass of silver nitrate using ration:

m(AgNO3) = n*M(AgNO3) where n is number of moles; M is molar mass in g/mol.

Thus:

m(AgNO3) = 1.23*(108 + 14 + 3*16) = 209.1 (g)

Using the chemical reaction we can determine the mass of silver chromate:

209.1 g X g

2AgNO3 + K2CrO4 = Ag2CrO4 (s) + 2KNO3

2*170 332

Make a proportion:

209.1/340 = X/332

where:

X = m (Ag2CrO4) = 209.1*332/340 = 204.18 (g)


Answer: m (Ag2CrO4) = 204.18 g.